A photon of 498nm was emitted from a silicon atom. Calculate the energy of all the atomic levels of silicon and show that the atomic levels are quantized.


1
Expert's answer
2021-08-23T02:57:05-0400

"E=\frac{hc}{λ}"

where:

  • "h=6.626 \u00d7 10^{\u221234}" J is Planck's constant
  •  "c=3\u00d710^8m\/s" is the speed of light

"E=\frac{6.626×10^{-34}×3×10^8}{498}= 3.99×10^{-28}"